A charge of 4.50x10^-8C is placed in a uniform electric field that is directed vertically upward with a magnitude of 5.00x10^4N/C. What work is done by the electrical force when the charge moves 0.450m to the right? 0.800m downward? 2.60m at an angle of 45° from the horizontal?

1
Expert's answer
2022-03-30T13:42:02-0400

(1)

Work

"W=Fdcos\theta"


"W=qEdcos\theta\\W=4.50\times10^{-8}\times5\times10^4\times0.450\times cos 90=0J"


"W=qEdcos\theta\\W=4.50\times10^{-8}\times5\times10^4\times0.800\times cos 0=1.8\times10^{-3}J"

"W=qEdcos\theta\\W=4.50\times10^{-8}\times5\times10^4\times2.60\times cos 45°=4.13\times10^{-3}J"